OpenStudy (anonymous):

lim [a^x sin(b/a^x)], when a>1 is = ? x->infinity

OpenStudy (gorv):

\[\lim_{x \rightarrow \infty} \frac{ a^x*\sin(\frac{ b }{ a^x } )}{ }\]

OpenStudy (gorv):

u knw sinx/x when limit x ....>0 ????

OpenStudy (gorv):

\[\lim_{x \rightarrow \infty} \frac{ \sin \frac{ b }{ a^x } }{ \frac{ 1 }{ a^x} }\]

OpenStudy (gorv):

multiply and divide by b

OpenStudy (gorv):

\[\lim_{x \rightarrow \infty} \frac{ b*\sin(\frac{ b }{ a^x }) }{ b*\frac{ 1 }{ a^x }}\]

OpenStudy (gorv):

=b

OpenStudy (anonymous):

Thank you very much

OpenStudy (gorv):

here 1/x is there so as x goex to infinity so it will 1/infinity=0 finally and comes to sin0/0 so we can apply lim x...>0 sinx/x

OpenStudy (anonymous):

Yes, I got it. Thanks once again.